A company wants to invest $180,000 in three accounts, one bearing 4.5%, another 6%, and another 8% interest. Twice as much will be invested at 4.5% as at 8%, and the total interest earned the first year needs to be $10,400. How much should be invested at each rate?
Answer by checkley71(8403)

2*.045X+.08X+(180000-3X).06=10400 OR .09X+.08X+10800-.18X=10400 OR
.17X+10800-.18X=10400 OR -.01X=-400 OR X=40000
PROOF .09*40000+.08*40000+(180000-3*40000).06=10400 OR
3600+3200+(180000-120000).06=10400 OR 6800+(60000*.06)=10400 OR
6800+3600=10400 OR 10400=10400